OpenStudy (anonymous):

Solve by completing the square. 3x^2-6x=2 Solve by completing the square. 3x^2-6x=2 @Mathematics

OpenStudy (lalaly):

\[x^2-2x-\frac{2}{3}=0\]\[(x-\frac{2}{2})^2 - (\frac{2}{2})^2 -\frac{2}{3} =0\]\[(x-1)^2-1-\frac{2}{3} =0\]\[(x-1)^2=\frac{5}{3}\]\[x-1 = \pm \sqrt{\frac{5}{3}}\]\[x=1 \pm \sqrt{\frac{5}{3}}\]

OpenStudy (anonymous):

Equation becomes : 3x^2-6x-2=0 here , a = 3 b = -6 c = -2 divide whole equation by a (3) Equation becomes = x^2 - 2x - 2/3 = 0 => x^2 - 2x = 2/3 Add (b/2a)^2 to both sides.. => x^2-2x +(3/2)^2 = 2/3 + (3/2)^2 => (x- 3/2)^2 = 2/3 + 9/4 => x = SQRT(2/3+9/4) + 3/2 Calculate it :)

OpenStudy (lalaly):

\[x^2 - bx +c =0\]completin the square \[(x-\frac{b}{2})^2 - (\frac{b}{2})^2 +c=0\] This is what i did ^

OpenStudy (anonymous):

Thank you for your help you are always awesome.

OpenStudy (lalaly):

ur welcome:D

OpenStudy (anonymous):

why can't you just use the quadratic formula, which is derived from completing the square?